# Querying Autonomous Vehicle Point Clouds: Enhanced by 3D Object Counting with CounterNet

### v1

@@ 5.3. Query Optimization by Model Selection
Autonomous vehicle point cloud frames vary significantly, making object counting challenging for models with fixed partitions and overlap ratios. Small, numerous objects benefit from more partitions, while larger objects require fewer. We propose a model selection method that chooses the nearest model based on feature distance. To mitigate errors in model center estimation, we apply the Chernoff bound to adjust selection probabilistically. The algorithm can be found in Appendix 2.
@@
To quantify the probability of deviation, we assume point cloud frames are independent and apply the Chernoff bound (Hellman and Raviv, 1970), which provides an upper bound on the probability that $\hat{\omega}_{m_{j}}$ deviates from $\mu$ by more than a given margin $\epsilon$: $P(|\hat{\omega}_{m_{j}}-\mu|>\epsilon)\leq exp(-\frac{n\epsilon^{2}}{2\xi^{2}})$.
Here, $\epsilon$ is the expected deviation, $\xi$ is the sample variance, and $n$ is the number of samples. In this equation, $\epsilon$ is the distance threshold defined by the user. Lower $\epsilon$ provides a higher probability. This probability $P_{m_{j}}$ quantifies our confidence in the estimated center $\hat{\omega}_{m_{j}}$, guiding adjustments to improve model selection.
Find the optimal model. During validation, for each new frame $p$, we compute the distance between $p$ and the center of each model $m\in M$, . The model with the shortest distance to the new point cloud frame is selected for further processing. Additionally, when selecting the closest model center, we must take into account the potential deviation of each model’s estimated center from its true center. For the distance between $p$ to each $m\in M$, we combine the distance with the probability: $d_{adjusted}(p,m_{j})=d(p,\hat{\omega}_{m_{j}})\times P_{m_{j}}$, where $d(p,\hat{\omega}_{m_{j}})$ is the distance between $p$ and $\hat{\omega}_{m_{j}}$, and $P_{m_{j}}$ is the confidence probability derived from the Chernoff bound. By incorporating $P_{m_{j}}$, we adjust for potential deviation in the estimated center and improve robustness in model selection.
@@ 6.1. Experiment Setup
All experiments were conducted on a server running Ubuntu 22.04.1, with Intel(R) Xeon(R) CPU E5-2697A v4 @ 2.60GHz and RTX3080. The code is implemented with OpenPCDet Framework (Team, 2020). Our code is available at (Authors, 2025).
Datasets. Experiments are conducted on the nuScenes (Caesar et al., 2019), KITTI (Behley et al., 2019) and Waymo-partial (Sun et al., 2020) (detailed explanation in Appendix G).
@@
Table 1. Notations for all tables and figures
@@
Combinations VN TF $CN$ $CN_{pt}$ $CN_{pt}^{o}$
(car, pedestrian) 0.582 0.425 0.700 0.800 0.842
(car, barrier) 0.616 0.506 0.693 0.808 0.852
@@
0.474 0.372 0.623 0.724 0.759
(truck, bus) 0.725 0.605 0.771 0.758 0.828
(bus, trailer) 0.844 0.890 0.910 0.897 0.913
(truck, trailer) 0.732 0.607 0.791 0.773 0.832
@@
COUNT
nuScenes KITTI Waymo
baselines VN 0.658 0.763
TF 0.683
CP 0.771
ours $CN$ 0.759 0.7959 0.779
$CN_{pt}$ 0.791 0.780 0.771
$CN_{pt}^{o}$ 0.857 0.846 0.807
Table 6. Overall query result of COUNT
Methods for Comparison. As explained in Section 2, unlike prior querying studies that focus on indexing (Li et al., 2025; Bang et al., 2023; Kang et al., 2019, 2022), our solution focuses on facilitating querying performance by improving object counting. As a result, we use the following detection models as our baselines for querying.
(1) VoxelNeXt (Chen et al., 2023) (VN): We use VoxelNeXt, which relies on sparse voxel features to facilitate detection.
(2) TransFusion (Bai et al., 2022) (TF): TransFusion utilizes an attention mechanism to fuse point cloud and image data for improved detection.
(3) CenterPoint (Yin et al., 2021) (CP): CenterPoint is the pioneer study of the center-based detection model.
Given the availability of baseline implementations on different datasets, we use VoxelNeXt and TransFusion as baselines for the nuScenes dataset. For the Waymo dataset, we adopt VoxelNeXt, while for the KITTI dataset, we use CenterPoint as the baseline.
@@
(1) RETRIEVAL. We experimented with two types of retrieval queries: RETRIEVAL-SINGLE and RETRIEVAL-MULTIPLE. These queries retrieve frames containing a specific number of objects. In most analytical scenarios—such as congestion detection—precise object counts are not strictly necessary; instead, a degree of deviation is acceptable to account for natural variability. To evaluate performance, we calculated the percentage of frames in which the retrieval condition was correctly met, allowing for 10% and 5% error tolerance. Since the number of objects in a frame can vary significantly (e.g., some frames contain over 50 cars, while others have at most 20 trucks, as shown in Table 11), we set the error tolerance per category proportionally. Specifically, the tolerance threshold for each object category is determined by multiplying the maximum number of that object type in a scene by the error tolerance. This approach ensures a fair evaluation across categories.
@@
(3) AGGREGATION. The Aggregation query provides statistical information, such as the total number of objects in a set of frames. We evaluate this using the SUM() operation. For random consecutive frame groups of lengths between 100 and 500, we calculate the sum of the queried objects and compare the evaluated sum to the ground truth across 1000 random queries. We use two metrics to evaluate the results: absolute difference and Q-error. The absolute difference measures the average deviation between the predicted and ground truth sums, providing a straightforward sense of numerical accuracy. In contrast, Q-error (quantile error) evaluates the ratio between the predicted and actual values, capturing the relative error, which is used widely in cardinality analysis for databse (Han et al., 2021). Specifically, it is defined as the maximum of the ratio between the prediction and the ground truth, and vice versa, i.e., $Q\text{-error}=\max\left(\frac{\hat{y}}{y},\frac{y}{\hat{y}}\right)$. Within the 1000 queries, we take the median as our reported Q-error.
@@ 6.2. Effectiveness Study on Various Query Types
Table 2 and Table 6 demonstrate the overall performance of the three query types: RETREIVAL, AGGREGATION, and COUNT. Across all three queries, the partition-based model ($CN_{pt}$) does not consistently outperform the baseline ($CN$), with performance drops particularly obvious on the KITTI dataset in both RETRIEVAL-SINGLE (i.e., RETRIEVAL-S) and AGGREGATION (i.e. AGG) queries. This decline is likely due to KITTI’s smaller spatial coverage per point cloud frame, making it more sensitive to partitioning. However, introducing overlap in the partitioning strategy ($CN_{pt}^{o}$) effectively mitigates this issue, delivering the best performance in most cases and demonstrating greater robustness across datasets of varying scales. Although all models exhibit decreased RETRIEVAL performance under the stricter 5% error tolerance, the degradation is more pronounced in $CN_{pt}$, while $CN_{pt}^{o}$ remains comparatively stable, further highlighting its resilience.
Table 3 - Table 5 show RETRIEVAL-MULTIPLE results on different datasets. In table 3, we evaluate two groups based on occurrence frequency: high-frequency objects (cars, pedestrians, barriers) and low-frequency objects (trucks, buses, trailers), by referring to Table 11 for object frequency on nuScenes in Appendix E. (1) High-Frequency Objects: Partitions enhance the result, and overlap further improves it, except for pedestrians and barriers, where duplicate counts in overlapped regions reduce accuracy due to their smaller size and higher occurrence rates. (2) Low-Frequency Objects: Partitions do not improve the accuracy, but overlap significantly enhances performance, particularly for sparsely distributed or large objects. For the detailed categorical results and analysis can be found in Appendix E - G.
@@ 6.3. Parameter Study
(3) Increasing Overlap Ratio: As demonstrated in Figure 12, increasing the expansion ratio contributes little to the improvement of query accuracy. This finding indicates that while overlap helps, its benefits are not substantial by blind expansion.
@@ 6.5. Study of Model Selection
Table 8 shows model selection results based on data distribution (i.e., frame allocation) with $\epsilon$ values from 0.1 to 0.2, where higher $\epsilon$ indicates a tighter bound. Key observations: (1) Model $CN_{pt=4}^{o=0.2}$ dominates, but frame allocation varies with adjustment. Without adjustment, many frames (e.g., 1116) are inefficiently allocated to suboptimal model $CN_{pt=2}^{o=0}$, while adjustment reallocates frames to more reliable models like $CN_{pt=2}^{o=0.1}$ (better for larger objects like buses), $CN_{pt=4}^{o=0}$, and $CN_{pt=9}^{o=0.1}$ (better for smaller objects like pedestrians). (2) Lower $\epsilon$ increasing adjustment rates and true center deviations, while larger $\epsilon$ reduce adjustment influence; a balance is crucial. The selection of $\epsilon$ can be found in Appendix E.3 (3) Evaluation via object count accuracy (Figure 11) using $CN_{pt=4}^{o=0.2}$ as the baseline shows that, without model selection, performance worsens due to $CN_{pt=2}^{o=0}$. Adjustments improve balance across categories, trading $CN_{pt=4}^{o=0.2}$’s exceptional detection of cars and buses for gains in detecting pedestrians, bicycles, and barriers. For more experiments of the model selection, see Appendix E.3.
@@ 7. Conclusion
In this paper, we present CounterNet, a method to facilitate querying 3D point cloud data with a focus on accurate object counting. By emphasizing object appearance detection and introducing a partition-based strategy with overlapping regions, we improve performance across varying object densities. Additionally, our dynamic model selection ensures robustness across scenarios. Experiments show that our approach enhances object counting accuracy, thereby improving query effectiveness. In the future, this research can be extended in several directions: (1) Investigating more efficient solutions for querying point cloud data, optimizing both performance and scalability. (2) Leveraging the query tool to support data acquisition research, enabling the extraction of relevant data from a data pool to enhance model training for specific tasks.
Appendix A Table of Notation
@@
Table 9. Table of notation
@@ Appendix B Motivation Example for the Study
Querying specific frames in point cloud data is crucial for facilitating target tasks. For example, to enhance a model’s ability to detect pedestrians, a direct and effective approach is to train or fine-tune the model using frames containing pedestrians rather than relying on the entire dataset. This targeted approach not only improves performance but also reduces computation costs. We conducted a simple experiment using the KITTI dataset. As shown in Figure 10, we compared pedestrian detection performance under three conditions: (1)Training on the full dataset (3712 frames). (2)Training on a randomly sampled subset of 1,000 frames. (3)Training on a carefully selected subset of frames containing pedestrians (about 955 frames).
The results demonstrated that using the carefully selected subset improved pedestrian detection performance, achieving similar or better results compared to the full dataset while reducing training costs by approximately two-thirds.
@@ C.1. Data Model
As illustrated below, our data model is designed with the NoSQL format, which brought more flexibility in aggregating data from different sources, such as different vehicles, making it more suitable for handling the diverse and complex nature of autonomous driving data. Object counting, being one of the most essential pieces of information required for querying (Kang et al., 2022, 2019; Cao et al., 2022; Bang et al., 2023), shapes the design of our data model, as shown below. We store key information from each frame, including object type, count, and the approximate positions of objects in a bird’s eye view, ensuring that the data model is optimized for efficient querying.
@@ Appendix D Methodology
D.1. Overlapped Partition
@@ 22  $centers=combineCenters(centers,r)$;
D.2. Model Selection
The overall process of model selection is explained in Algorithm 2, which main contains three steps: (1) Model center representation. (2) Center adjustment with confidence. (3) Find the optimal model.
@@ E.2. Categorical Query Results
By analyzing the results from RETRIEVAL-SINGLE (Table 12), COUNT and AGGREGATION (Table 13), we observe: (1) CounterNet (CN) Performance: CounterNet outperforms detection models like VoxelNeXt (VN) and TransFusion (TF), demonstrating its outstanding in extracting query-specific information. (2) Impact of Partitioning: Partitioning improves accuracy for small-object categories like cars and barriers, while performance remains stable for pedestrians. However, it reduces accuracy for larger objects (e.g., trucks, buses) due to increased segmentation across partitions, aligning with our hypothesis that partitioning benefits small or numerous objects but hinders performance for infrequent, large objects. (3) Impact of Overlapping: Overlapped partitions improve performance across all categories, mitigating partitioning drawbacks (e.g., segmentation issues) while maintaining its advantages for counting small or numerous objects.
@@
VN TF $CN$ $CN_{pt}$ $CN_{pt}^{o}$ VN TF $CN$ $CN_{pt}$ $CN_{pt}^{o}$
car 0.801 0.615 0.807 0.919 0.967 0.536 0.506 0.601 0.774 0.861
truck 0.778 0.621 0.802 0.799 0.858 0.594 0.610 0.659 0.609 0.705
bus 0.917 0.959 0.957 0.944 0.960 0.917 0.959 0.957 0.944 0.960
trailer 0.915 0.925 0.953 0.950 0.951 0.815 0.909 0.912 0.880 0.871
const 0.845 0.886 0.929 0.910 0.931 0.849 0.886 0.929 0.909 0.810
ped 0.711 0.640 0.870 0.869 0.873 0.469 0.553 0.741 0.735 0.748
moto 0.728 0.867 0.933 0.886 0.940 0.728 0.866 0.933 0.886 0.940
bicyc 0.632 0.828 0.940 0.898 0.942 0.632 0.907 0.940 0.898 0.942
cone 0.706 0.722 0.933 0.926 0.954 0.528 0.717 0.875 0.874 0.867
barrier 0.761 0.802 0.874 0.883 0.883 0.640 0.755 0.811 0.823 0.824
@@
VN TF $CN$ $CN_{pt}$ $CN_{pt}^{o}$ VN TF $CN$ $CN_{pt}$ $CN_{pt}^{o}$ VN TF $CN$ $CN_{pt}$ $CN_{pt}^{o}$
car 0.552 0.355 0.947 0.927 0.953 4.021 6.488 3.986 2.357 0.662 1.478 1.736 1.401 1.403 1.290
truck 0.309 0.431 0.351 0.588 0.592 1.544 2.568 1.157 0.723 0.506 1.979 2.222 1.828 1.842 1.522
bus 0.950 0.994 0.859 0.857 0.860 0.401 0.220 0.228 0.259 0.183 1.000 1.000 1.113 1.065 1.000
trailer 0.562 0.858 0.787 0.829 0.865 0.786 0.703 0.373 0.355 0.390 1.001 1.256 1.057 1.144 1.240
const 0.920 0.947 0.890 0.851 0.936 0.628 0.475 0.221 0.235 0.275 1.220 1.180 1.112 1.131 1.087
ped 0.262 0.307 0.671 0.732 0.745 6.039 7.537 2.768 2.636 2.516 2.560 2.913 2.157 2.143 1.967
moto 0.866 0.920 0.825 0.830 0.884 1.021 0.573 0.190 0.304 0.201 1.572 1.232 1.315 1.293 1.061
bicyc 0.771 0.946 0.949 0.948 0.950 1.315 0.649 0.144 0.330 0.140 1.764 1.482 1.167 1.177 1.064
cone 0.180 0.251 0.572 0.586 0.892 5.189 4.854 1.091 1.123 0.898 3.161 2.725 1.550 1.598 1.944
barrier 0.208 0.410 0.745 0.765 0.896 5.429 4.029 2.128 2.115 1.670 2.303 1.996 1.464 1.490 1.041

### v2

@@ 5.3. Query Optimization by Model Selection
Autonomous vehicle point cloud frames vary significantly, making object counting challenging for models with fixed partitions and overlap ratios. Small, numerous objects benefit from more partitions, while larger objects require fewer. We propose a model selection method that chooses the nearest model based on feature distance. To mitigate errors in model center estimation, we apply the Chernoff bound to adjust selection probabilistically. The algorithm can be found in Appendix E.3.
@@
To quantify the probability of deviation, we assume point cloud frames are independent and apply the Chernoff bound (Hellman and Raviv, 1970), which provides an upper bound on the probability that $\hat{\omega}_{m_{j}}$ deviates from $\mu$ by more than a given margin $\epsilon$: $P(|\hat{\omega}_{m_{j}}-\mu|>\epsilon)\leq exp(-\frac{n\epsilon^{2}}{2\xi^{2}})$. Here, $\epsilon$ is the expected deviation, $\xi$ is the sample variance, and $n$ is the number of samples. In this equation, $\epsilon$ is the distance threshold defined by the user. Lower $\epsilon$ provides a higher probability. This probability $P_{m_{j}}$ quantifies our confidence in the estimated center $\hat{\omega}_{m_{j}}$, guiding adjustments to improve model selection.
Finding the optimal model. During validation, for each new frame $p$, we compute the distance between $p$ and the center of each model $m\in M$, . The model with the shortest distance to the new point cloud frame is selected for further processing. Additionally, when selecting the closest model center, we must take into account the potential deviation of each model’s estimated center from its true center. For the distance between $p$ to each $m\in M$, we combine the distance with the probability: $d_{adjusted}(p,m_{j})=d(p,\hat{\omega}_{m_{j}})\times P_{m_{j}}$, where $d(p,\hat{\omega}_{m_{j}})$ is the distance between $p$ and $\hat{\omega}_{m_{j}}$, and $P_{m_{j}}$ is the confidence probability derived from the Chernoff bound. By incorporating $P_{m_{j}}$, we adjust for potential deviation in the estimated center and improve robustness in model selection.
@@ 6.1. Experiment Setup
All experiments were conducted on a server running Ubuntu 22.04.1, with Intel(R) Xeon(R) CPU E5-2697A v4 @ 2.60GHz and RTX3080. The code was implemented with OpenPCDet Framework (Team, 2020). Our code is available at (Authors, 2025).
Datasets. Experiments were conducted on the nuScenes (Caesar et al., 2019), KITTI (Behley et al., 2019) and Waymo-partial (Sun et al., 2020) (detailed explanation in Appendix G).
@@
Table 1. Notations in experiments
@@
Combinations VN TF SN $CN$ $CN_{pt}$ $CN_{pt}^{o}$
(car, pedestrian) 0.582 0.425 0.675 0.700 0.800 0.842
(car, barrier) 0.616 0.506 0.698 0.693 0.808 0.852
@@
0.474 0.372 0.571 0.623 0.724 0.759
(truck, bus) 0.725 0.605 0.794 0.771 0.758 0.828
(bus, trailer) 0.844 0.890 0.921 0.910 0.897 0.913
(truck, trailer) 0.732 0.607 0.784 0.791 0.773 0.832
@@
Methods for Comparison. As explained in Section 2, unlike prior querying studies that focus on indexing (Li et al., 2025; Bang et al., 2023; Kang et al., 2019, 2022), our solution focuses on facilitating querying performance by improving object counting. As a result, we use the following detection models as our baselines for querying: (1) VoxelNeXt (Chen et al., 2023) (VN), which relies on sparse voxel features to facilitate detection; (2) TransFusion (Bai et al., 2022) (TF), which utilizes an attention mechanism to fuse point cloud and image data for improved detection; (3) CenterPoint (Yin et al., 2021) (CP), a pioneer study of the center-based detection model; (4) SAFDNet (Zhang et al., 2024) (SN), which utilises a fully sparse feature map for object detection. SAFDNet is one of the most recent detection studies, and shows better performance than other baselines.
Given the availability of baseline implementations on different datasets, we use VoxelNeXt, TransFusion and SAFDNet as baselines for the nuScenes dataset. For the Waymo dataset, we adopt VoxelNeXt, while for the KITTI dataset, we adopt CenterPoint.
@@
(1) RETRIEVAL. We experiment with two types of retrieval queries: RETRIEVAL-SINGLE and RETRIEVAL-MULTIPLE. These queries retrieve frames containing a specific number of objects. In most analytical scenarios, such as congestion detection, precise object counts are not strictly necessary; instead, a degree of deviation is acceptable to account for natural variability. To evaluate performance, we calculate the percentage of frames in which the retrieval condition was correctly met, allowing for 10% and 5% error tolerance. Specifically, the tolerance threshold for each object category is determined by multiplying the maximum number of that object type in a scene by the error tolerance. This approach ensures a fair evaluation across categories.
@@
(3) AGGREGATION. The aggregation query provides statistical information such as total number of objects in a set of frames. We evaluate this using the SUM() operation. For random consecutive frame groups of lengths between 100 and 500, we calculate the sum of queried objects and compare it to the ground truth across 1000 random queries. We use two metrics to evaluate the results: absolute difference and Q-error. The absolute difference measures the average deviation between the predicted and ground truth sums, providing a straightforward sense of numerical accuracy. In contrast, Q-error (quantile error) evaluates the ratio between the predicted and actual values, capturing the relative error (Han et al., 2021).
@@ 6.2. Effectiveness Study on Various Query Types
Table 2 and Table 6 demonstrate the overall performance of the three query types: RETRIEVAL, AGGREGATION, and COUNT. Across all three queries, the partition-based model ($CN_{pt}$) does not consistently outperform the baseline ($CN$), with performance drops particularly obvious on the KITTI dataset in both RETRIEVAL-SINGLE (i.e., RETRIEVAL-S) and AGGREGATION (i.e. AGG) queries. This decline is likely due to KITTI’s smaller spatial coverage per point cloud frame, making it more sensitive to partitioning. However, introducing overlap in the partitioning strategy ($CN_{pt}^{o}$) effectively mitigates this issue, delivering the best performance in most cases and demonstrating greater robustness across datasets of varying scales. Although all models exhibit decreased RETRIEVAL performance under the stricter 5% error tolerance, the degradation is more pronounced in $CN_{pt}$, while $CN_{pt}^{o}$ remains comparatively stable, further highlighting its resilience.
Tables 3-5 show RETRIEVAL-MULTIPLE results on different datasets. In Table 3, we evaluate two groups based on occurrence frequency: high-frequency objects (cars, pedestrians, barriers) and low-frequency objects (trucks, buses, trailers), by referring to Table 11 for object frequency on nuScenes in Appendix E. (1) High-Frequency Objects: Partitions enhance the result, and overlap further improves it, except for pedestrians and barriers, where duplicate counts in overlapped regions reduce accuracy due to their smaller size and higher occurrence rates. (2) Low-Frequency Objects: Partitions do not improve the accuracy, but overlap significantly enhances performance, particularly for sparsely distributed or large objects. The detailed categorical results and analysis can be found in Appendix E - G.
@@ 6.3. Parameter Study
(3) Increasing Overlap Ratio: As demonstrated in Figure 12, increasing the expansion ratio contributes little to the improvement of query accuracy. This finding indicates that while overlap helps, its benefits are not substantial through blind expansion.
@@ 6.4. Efficiency Study
COUNT
nuScenes KITTI Waymo
baselines VN 0.658 0.763
TF 0.683
SN 0.718
CP 0.771
ours $CN$ 0.759 0.7959 0.779
$CN_{pt}$ 0.791 0.780 0.771
$CN_{pt}^{o}$ 0.857 0.846 0.807
Table 6. Overall query result of COUNT
6.5. Study of Model Selection
@@
Table 8 shows model selection results based on data distribution (i.e., frame allocation) with $\epsilon$ values from 0.1 to 0.2, where higher $\epsilon$ indicates a tighter bound. Key observations: (1) Model $CN_{pt=4}^{o=0.2}$ dominates, but frame allocation varies with adjustment. Without adjustment, many frames (e.g., 1116) are inefficiently allocated to suboptimal model $CN_{pt=2}^{o=0}$, while adjustment reallocates frames to more reliable models like $CN_{pt=2}^{o=0.1}$ (better for larger objects like buses), $CN_{pt=4}^{o=0}$, and $CN_{pt=9}^{o=0.1}$ (better for smaller objects like pedestrians). (2) Lower $\epsilon$ increases adjustment rates and true center deviations, while larger $\epsilon$ reduce adjustment influence; a balance is crucial. The selection of $\epsilon$ can be found in Appendix E.3. (3) Evaluation via object count accuracy (Figure 11) using $CN_{pt=4}^{o=0.2}$ as the baseline shows that, without model selection, performance worsens due to $CN_{pt=2}^{o=0}$. Adjustments improve balance across categories, trading $CN_{pt=4}^{o=0.2}$’s exceptional detection of cars and buses for gains in detecting pedestrians, bicycles, and barriers. For additional experiments on model selection, see Appendix E.3.
@@ 7. Conclusion
In this paper, we present CounterNet, a method to facilitate querying 3D point cloud data with a focus on accurate object counting. By emphasizing object appearance detection and introducing a partition-based strategy with overlapping regions, we improve performance across varying object densities. Additionally, our dynamic model selection ensures robustness across scenarios. In the future, this research can be extended in directions: (1) Investigating more efficient solutions for querying point cloud data, optimizing both performance and scalability. (2) Leveraging the query tool to support data acquisition research, enabling the extraction of relevant data from a data pool to enhance model training for specific tasks.
ACKNOWLEDGEMENT
Xiaoyu Zhang is a recipient of CSIRO Scholarship Stipend and Top-Up. This research was also supported partially by the Australian Government through the Australian Research Council’s Discovery Projects funding scheme (projects DP240101211 and DP220101823).
Appendix A Table of Notations
@@
Table 9. Table of notations
@@ Appendix B Motivation Example for the Study
Querying specific frames in point cloud data is crucial for facilitating target tasks. For example, to enhance a model’s ability to detect pedestrians, a direct and effective approach is to train or fine-tune the model using frames containing pedestrians rather than relying on the entire dataset. This targeted approach not only improves performance but also reduces computation costs. We conduct a simple experiment using the KITTI dataset. As shown in Figure 10, we compare pedestrian detection performance under three conditions: (1) Training on the full dataset (3712 frames). (2) Training on a randomly sampled subset of 1,000 frames. (3) Training on a carefully selected subset of frames containing pedestrians (about 955 frames).
The results demonstrate that using the carefully selected subset improved pedestrian detection performance, achieving similar or better results compared to the full dataset while reducing training costs by approximately two-thirds.
@@ C.1. Data Model
As illustrated below, our data model is designed with the NoSQL format, which brings more flexibility in aggregating data from different sources, such as different vehicles, making it more suitable for handling the diverse and complex nature of autonomous driving data. Object counting, being one of the most essential pieces of information required for querying (Kang et al., 2022, 2019; Cao et al., 2022; Bang et al., 2023), shapes the design of our data model, as shown below. We store key information from each frame, including object type, count, and the approximate positions of objects in a bird’s eye view, ensuring that the data model is optimized for efficient querying.
@@ Appendix D Methodology
D.1. Steps of Feature Map Partition with Overlap
@@ 22  $centers=combineCenters(centers,r)$;
D.2. Steps of Model Selection
The overall process of model selection is explained in Algorithm 2, which contains three main steps: (1) Model center representation. (2) Center adjustment with confidence. (3) Find the optimal model.
@@ E.2. Categorical Query Results
By analyzing the results from RETRIEVAL-SINGLE (Table 12), COUNT and AGGREGATION (Table 13), we observe: (1) CounterNet (CN) Performance: CounterNet outperforms detection models, demonstrating its outstanding performance in extracting query-specific information. (2) Impact of Partitioning: Partitioning improves accuracy for small-object categories like cars and barriers, while performance remains stable for pedestrians. However, it reduces accuracy for larger objects (e.g., trucks, buses) due to increased segmentation across partitions, aligning with our hypothesis that partitioning benefits small or numerous objects but hinders performance for infrequent, large objects. (3) Impact of Overlapping: Overlapped partitions improve performance across all categories, mitigating partitioning drawbacks (e.g., segmentation issues) while maintaining their advantages for counting small or numerous objects.
@@
VN TF SN $CN$ $CN_{pt}$ $CN_{pt}^{o}$ VN TF SN $CN$ $CN_{pt}$ $CN_{pt}^{o}$
car 0.801 0.615 0.861 0.807 0.919 0.967 0.536 0.506 0.635 0.601 0.774 0.861
truck 0.778 0.621 0.805 0.802 0.799 0.858 0.594 0.610 0.668 0.659 0.609 0.705
bus 0.917 0.959 0.982 0.957 0.944 0.960 0.917 0.959 0.982 0.957 0.944 0.960
trailer 0.915 0.925 0.937 0.953 0.950 0.951 0.815 0.909 0.890 0.912 0.880 0.871
const 0.845 0.886 0.924 0.929 0.910 0.931 0.849 0.886 0.924 0.929 0.909 0.810
ped 0.711 0.640 0.773 0.870 0.869 0.873 0.469 0.553 0.579 0.741 0.735 0.748
moto 0.728 0.867 0.906 0.933 0.886 0.940 0.728 0.866 0.906 0.933 0.886 0.940
bicyc 0.632 0.828 0.893 0.940 0.898 0.942 0.632 0.907 0.893 0.940 0.898 0.942
cone 0.706 0.722 0.824 0.933 0.926 0.954 0.528 0.717 0.727 0.875 0.874 0.867
barrier 0.761 0.802 0.810 0.874 0.883 0.883 0.640 0.755 0.738 0.811 0.823 0.824
@@
VN TF SN $CN$ $CN_{pt}$ $CN_{pt}^{o}$ VN TF SN $CN$ $CN_{pt}$ $CN_{pt}^{o}$ VN TF SN $CN$ $CN_{pt}$ $CN_{pt}^{o}$
car 0.552 0.355 0.695 0.947 0.927 0.953 4.021 6.488 3.213 3.986 2.357 0.662 1.478 1.736 1.634 1.401 1.403 1.290
truck 0.309 0.431 0.602 0.351 0.588 0.592 1.544 2.568 1.385 1.157 0.723 0.506 1.979 2.222 1.983 1.828 1.842 1.522
bus 0.950 0.994 0.995 0.859 0.857 0.860 0.401 0.220 0.112 0.228 0.259 0.183 1.000 1.000 1.292 1.113 1.065 1.000
trailer 0.562 0.858 0.798 0.787 0.829 0.865 0.786 0.703 0.588 0.373 0.355 0.390 1.001 1.256 1.075 1.057 1.144 1.240
const 0.920 0.947 0.904 0.890 0.851 0.936 0.628 0.475 0.287 0.221 0.235 0.275 1.220 1.180 1.166 1.112 1.131 1.087
ped 0.262 0.307 0.425 0.671 0.732 0.745 6.039 7.537 4.726 2.768 2.636 2.516 2.560 2.913 2.023 2.157 2.143 1.967
moto 0.866 0.920 0.887 0.825 0.830 0.884 1.021 0.573 0.406 0.190 0.304 0.201 1.572 1.232 1.272 1.315 1.293 1.061
bicyc 0.771 0.946 0.956 0.949 0.948 0.950 1.315 0.649 0.437 0.144 0.330 0.140 1.764 1.482 1.201 1.167 1.177 1.064
cone 0.180 0.251 0.433 0.572 0.586 0.892 5.189 4.854 3.055 1.091 1.123 0.898 3.161 2.725 2.288 1.550 1.598 1.944
barrier 0.208 0.410 0.494 0.745 0.765 0.896 5.429 4.029 3.574 2.128 2.115 1.670 2.303 1.996 1.621 1.464 1.490 1.041